OpenStudy (anonymous):

find all solution of the given equation in the interval [0,2pi) tan(x/2)=sinx

OpenStudy (praxer):

@ mae23 here, we can write tan(x/2) as sin(x/2)/cos(x/2) then , $$\huge \dfrac{\sin(x/2)}{\cos(x/2}=sinx$$ now, we know that sinx/2 can be written as 2sin(x/2)cos(x/2) than $$\huge \dfrac{\sin(x/2)}{\cos(x/2}=2sin(x/2)cos(x/2)$$ now, $$\huge 1=2\cos^2(x/2)$$ finally, $$\huge \cos(x/2)=\dfrac {1}{\sqrt{2}}$$ we know that $$\huge \cos(\pi/4)= 1/\sqrt{2}$$ so,$$\huge x/2= \pi/4 $$ hence, $$\huge x= \pi/2$$ Understood it :)

OpenStudy (praxer):

@mae23

OpenStudy (praxer):

$$ \huge \color{red} {\ sinx = 2\ sin(x/2)\ cos(x/2)}$$

OpenStudy (anonymous):

yes . thank you

OpenStudy (praxer):

$$\huge\frak \color{red} {WELCOME........}$$

OpenStudy (anonymous):

@praxer is the interval equation pi/2,pi?

OpenStudy (praxer):

no it is not.

OpenStudy (praxer):

u gave the interval to be $$\huge \color{red} {{{[0,\pi}]}}$$

OpenStudy (praxer):

sorry [0,2pi]

OpenStudy (praxer):

it is in [0,2pi]

OpenStudy (anonymous):

oh? I taught 2pi is not included

OpenStudy (praxer):

see $$\huge \pi/2$$ lies in the interval you denoted.